# Unitarity, Causality, and Solar System Bounds, Significantly Limit Using Gravitational Waves to Test General Relativity

### v3

@@ Abstract
The prospect of detecting/constraining deviations from general relativity by studying gravitational waves (GWs) from merging black holes has been one of the primary motivations of GW interferometers like LIGO/Virgo. Within pure gravity, the only possible way deviations can arise is from the existence of higher order derivative corrections, namely higher powers of the Riemann curvature tensor, in the effective action. Any observational bounds imply constraints on the corresponding Wilson coefficients. At the level of the action, one can imagine the coefficients are sufficiently large so as to be in principle detectable. However, from the point of view of some fundamental principles, namely causality and unitarity, this is much less clear, as we examine here. We begin by reviewing certain known bounds on these coefficients, which together imply a low cut off on the effective theory. We then consider a possible mechanism to generate such terms, namely in the form of many minimally coupled light scalars that can be integrated out to give these higher order operators. We show that a by product of this is the generation of quantum corrections to Newton’s potential, whose observable consequences are already ruled out by solar system tests. We point out that over 7 orders of magnitude of improvement in interferometer sensitivity would be required to avoid such solar system constraints.
@@ 1 Introduction
Searches for either new effects or corrections that deviate from general relativity (GR) is an ongoing front in both the experimental and theoretical aspects of gravitation. On the experimental side, we have a century of precision tests of GR in the solar system. However, within the solar system, the spatial curvature is very small. More recently, we have LIGO/Virgo LIGOScientific:2016aoc detecting gravitational waves from merging black holes, where the spatial curvature is larger. Currently, all observations are broadly compatible with the predictions of GR. On the theoretical side, much has been devoted to understanding the space of possible deviations. Within effective field theory (EFT), the only way deviations can appear is through higher derivative operators and their possible appearance within an ultra-violated (UV) completion.
@@
In Section 2 we lay out the higher derivative operators within the EFT for gravity to discuss the causality bounds on the Wilson coefficients. This includes cubic operators (with coefficients $\Lambda_{3}$) and quartic operators (with coefficients $\Lambda_{4}$). We generate these higher derivative operators from a set of massive scalars that can be integrate out. We summarize bounds on the set of masses $m_{i}$ and the number of such particles $N$ can appear via loop corrections.
@@ 2.1 Building the Effective Field Theory
The effective field theory action for gravity is an expansion in derivatives. We set $\hbar=c=1$ with metric signature (+,,,)(+,-,-,-)( + , - , - , - ) (and adapting the notation of Caron-Huot:2022ugt ) we write the expansion as
$S_{\text{gravity}}=\int d^{4}x\sqrt{-g}\biggm{(}$ $-\Lambda-\frac{2R}{\kappa^{2}}+\frac{\alpha_{2}}{\kappa^{2}}R^{(2)}-\frac{1}{3\kappa^{2}}\left(\alpha_{3}R^{(3)}+\tilde{\alpha}_{3}\tilde{R}^{(3)}\right)$
@@
We have only included terms that involve powers of Riemann, but not terms that are powers of Ricci or Ricci tensor, such as $R^{2}$ or $R^{3}$, etc. The reason being is that we are primarily interested in the merger of Kerr black holes. Such black holes have $R=0$ and $R_{\mu\nu}=0$ as they are vacuum solutions. So we only need to focus on powers of Riemann, as this is non-zero even for vacuum solutions. One could consider these as local terms, $R^{2},\,R_{\mu\nu}R^{\mu\nu}$, etc, and inquire about their possible corrections to Newton’s law. We comment on this issue in Appendix C.
@@ 2.2 Constraints from Observed Gravitational Waves
In Ref. Sennett:2019bpc , bounds are placed on these operators from comparison to LIGO/Virgo data (although in Ref. Sennett:2019bpc the cubic term was assumed small, for reasons we will come to below). We can express these bounds in terms of a physical scales (inverse lengths) $\Lambda_{3}$ and $\Lambda_{4}$ that sets the size of the cubic and quartic powers of the Riemann operator in the action. These are related to our Wilson coefficients as
@@
$\Lambda^{-1}_{4}\lesssim\Lambda^{-1}_{4,0}\simeq 150\,\text{km}.$ (4)
(i.e., we are using the notation $\Lambda_{4,0}$ as short-handed notation for the inverse scale of 150 km). This scale naturally emerges since black holes of mass $\sim 30\,M_{sun}$ have a Schwarzschild radius of $\approx 90$ km, which is of this order. Although the cubic terms were not directly included in the analysis, a similar bound is expected for them also, i.e., $\Lambda_{3}^{-1}\lesssim\Lambda^{-1}_{3,0}\sim 150\,\text{km}$.
It is important to note that such a bound is stronger than current bounds from the solar system. The reason for this is as follows: even though solar system bounds have more precision, they involve much weaker curvatures. Furthermore, if we go to scales below this; for example down to several micro-meters, the Cavendish type tests of gravity are still less stringent, as again the curvatures produced by laboratory masses are so tiny. So at the level of classical field theory, these bounds from GWs of merging black holes provide the tightest known direct constraints on the coefficients of these higher dimension operators.
@@ 2.3 Constraints from Causality
As has been known for the past few years Camanho:2014apa , if one only includes $(\text{Riem})^{3}$ in the effective action for gravitation, and no other higher dimension operators, then there is superluminality. However, when the quartic operators $(\text{Riem})^{4}$ are included, then non-zero cubic terms are allowed, so long as an inequality between the cubic and quartic coefficients (which will be explained more in-depth in section 2.4), is satisfied (see Bern:2021ppb and Caron-Huot:2022ugt ). The inequality is
@@ 2.4 Higher Derivative Operators from Integrating Out Massive Particles
The above higher derivative operators ultimately arise from whatever is the underlying UV completion of gravity. However, since these Wilson coefficients flow with scale, we do not need to know the full UV completion details, such as string theory. Instead, we can imagine that there are many massive, but somewhat light, particles that we integrate out and we generate these operators in the infrared. As a concrete example, we shall focus on integrating out $N$ minimally coupled massive scalars $\phi_{i}$. The starting action is
$S=\int d^{4}x\sqrt{-g}\biggm{(}$ $-\Lambda-\frac{2R}{\kappa^{2}}+\sum_{i}\frac{1}{2}(\partial\phi_{i})^{2}-\sum_{i}\frac{1}{2}m_{i}^{2}\phi_{i}^{2}+\mathcal{L}_{\text{matter}}+\cdots\biggm{)}$ (9)
@@
[…]
@@
The numerical prefactors $(75/11)^{1/6}$ and $(15/4)^{1/4}$ are both approximately 1.4. Then if there is only one type of massive particle being integrated out this becomes
$\Lambda_{4}\geq 1.4\,m_{\text{min}},\quad\quad\Lambda_{3}\geq 1.4\,m_{\text{min}}.$ (14)
On the other hand, if there are several species of particles with different masses, then it is simple to see that the factor involving sums of powers of masses obeys
$\frac{\left(\sum_{i}(m_{\text{min}}/m_{i})^{2}\right)}{\left(\sum_{i}(m_{\text{min}}/m_{i})^{4}\right)}\geq 1$ (15)
since there is a fourth power in the denominator and only a second power in the numerator of terms that are each $m_{\text{min}}/m_{i}\leq 1$ from the definition of $m_{\text{min}}$. Here equality only occurs when all the masses are equal, but otherwise this factor is larger than unity. In the latter case, the above inequalities (13) are even sharper, i.e., the energy scales controlling the higher order operators in the EFT must be even larger than $1.4\,m_{\text{min}}$ the lightest mass being integrated out. Hence if the masses being integrated out are somewhat heavy, then these scales are correspondingly microscopic and one cannot use a large $N$ to alter this conclusion.
We can invert this logic and use it to bound $m_{\text{min}}$. The best case scenario for phenomenological purposes then is that all the masses are equal $m=m_{\text{min}}=m_{1}=\ldots=m_{N}$. If we use the observational bound on the EFT of Ref. Sennett:2019bpc of $\Lambda_{4,0}\simeq(150\text{km})^{-1}$ as a useful reference, then we obtain a bound on the mass of the lightest particle of
$m\leq m_{c}=9.5\times 10^{-13}\,\text{eV}(\Lambda_{4}/\Lambda_{4,0})$ (16)
@@
with equality holding when we saturate the theoretical bound $m=m_{c}$. Although this is an extremely large $N$, it is important to note that such species are assumed to be in a hidden sector and not directly coupled to the Standard Model at tree level. Their only direct couplings to the Standard Model occur via quantum loops as we now explore and learn their possible observational consequences. One might expect that such a huge number of particles would already have been easily visible at colliders, for example from $e^{+}e^{-}$ annihilation into these scalars via the graviton in the s-channel; but it is important to recall that the proposal that is presented in Ref. Sennett:2019bpc is that of the “soft UV completion", meaning that one should not use the theory at arbitrarily low scales.
@@ 3.1 Massive Scalar Amplitude
$\Pi_{\lambda\xi\mu\nu,1}(q)$ $=-\frac{1}{2}\mu^{4-d}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{\tau_{\lambda\xi}^{(1)}(l+q,l;m)\tau_{\mu\nu}^{(1)}(l,l+q;m)}{(l^{2}-m^{2})((l+q)^{2}-m^{2})}$
$=-\frac{1}{2}\mu^{4-d}\int\frac{d^{d}p}{(2\pi)^{d}}\frac{\tau_{\lambda\xi}^{(1)}(p+q,p;m)\tau_{\mu\nu}^{(1)}(p,p+q;m)}{(p^{2}-\Delta)^{2}}$ (23)
where $\Delta=m^{2}-x(1-x)q^{2}$ and $\mu$ is the scale at which we define the couplings, and we have suppressed the $+i0^{+}$ in the Feynman propagator for the sake of brevity. The 1PI changes for equation (3.1) as
$\Pi_{\lambda\xi\mu\nu,2}(q)=\mu^{4-d}i\int\frac{d^{d}l}{(2\pi)^{d}}\frac{\tau^{(2)}_{\lambda\xi\mu\nu}{(l;m)}}{l^{2}-m^{2}}$ (24)
@@ 3.1.1 One-Loop Diagrams
$A_{0}(m^{2})=\frac{\mu^{4-d}}{(2\pi)^{d}}\int\frac{d^{d}p}{i\pi}\frac{1}{p^{2}-m^{2}}$ (27)
@@
$B_{0}(p^{2},m^{2},m^{2})=$ $\frac{\mu^{4-d}}{(2\pi)^{d}}\int\frac{d^{d}q}{i\pi^{2}}\frac{1}{(q^{2}-m^{2})((q+p)^{2}-m^{2})}$ (29)

### v5

# Unitarity, Causality, and Solar System Bounds May Significantly Limit the Use of Gravitational Waves to Test General Relativity

@@ Abstract
The prospect of detecting/constraining deviations from general relativity by studying gravitational waves (GWs) from merging black holes has been one of the primary motivations of GW interferometers like LIGO/Virgo. Within pure gravity, the only possible way deviations can arise is from the existence of higher order derivative corrections, namely higher powers of the Riemann curvature tensor, in the effective action. Any observational bounds imply constraints on the corresponding Wilson coefficients. At the level of the action, one can imagine the coefficients are sufficiently large so as to be in principle detectable. However, from the point of view of some fundamental principles, namely causality and unitarity, this is much less clear, as we examine here. We begin by reviewing certain known bounds on these coefficients, which together imply a low cut off on the effective theory. We then consider a possible mechanism to generate such terms, namely in the form of many scalars, minimally coupled to only gravity, that can be integrated out to give these higher order operators. We show that a by product of this is the generation of quantum corrections to Newton’s potential, whose observable consequences are already ruled out by solar system tests. We point out that over 7 orders of magnitude of improvement in interferometer sensitivity would be required to avoid such solar system constraints. We also mention further constraints from Hawking radiation from black holes.
@@ 1 Introduction
Searches for either new effects or corrections that deviate from general relativity (GR) is an ongoing front in both the experimental and theoretical aspects of gravitation. On the experimental side, we have a century of precision tests of GR in the solar system. However, within the solar system, the spatial curvature is very small. More recently, we have LIGO/Virgo LIGOScientific:2016aoc detecting gravitational waves from merging black holes, where the spatial curvature is larger. Currently, all observations are broadly compatible with the predictions of GR. On the theoretical side, much has been devoted to understanding the space of possible deviations. Within effective field theory (EFT), the only way deviations can appear is through higher derivative operators and their possible appearance within an ultraviolet completion.
@@
In Section 2 we lay out the higher derivative operators within the EFT for gravity to discuss the causality bounds on the Wilson coefficients. This includes cubic operators (with coefficients $\Lambda_{3}$) and quartic operators (with coefficients $\Lambda_{4}$). We generate these higher derivative operators from a set of massive scalars minimally coupled to gravity that can be integrated out. We summarize bounds on the set of masses $m_{i}$ and the number $N$ of particles that can appear via loop corrections.
@@ 2.1 Building the Effective Field Theory
The effective field theory action for gravity is an expansion in derivatives. We set $\hbar=c=1$ with metric signature (+,,,)(+,-,-,-) (and adapting the notation of Caron-Huot:2022ugt ) we write the expansion as
$S_{\text{gravity}}=\uprightintop\displaylimits d^{4}x\sqrt{-g}\biggm($ $-\Lambda-\frac{2R}{\kappa^{2}}+\frac{\alpha_{2}}{\kappa^{2}}R^{(2)}-\frac{1}{3\kappa^{2}}\left(\alpha_{3}R^{(3)}+\tilde{\alpha}_{3}\tilde{R}^{(3)}\right)$
@@
We have only included terms that involve powers of Riemann, but not terms that are powers of Ricci or Ricci tensor, such as $R^{2}$ or $R^{3}$, etc. The reason being is that we are primarily interested in the merger of Kerr black holes, whose gravitational waves have been detected. Such black holes have $R=0$ and $R_{\mu\nu}=0$ as they are vacuum solutions. So we only need to focus on powers of Riemann, as this is non-zero even for vacuum solutions. One could consider these as local terms, $R^{2},\,R_{\mu\nu}R^{\mu\nu}$, etc, and inquire about their possible corrections to Newton’s law. We comment on this issue in Appendix C.
@@ 2.2 Constraints from Observed Gravitational Waves
In Ref. Sennett:2019bpc , bounds are placed on these curvature operators from comparison to LIGO/Virgo data (although in Ref. Sennett:2019bpc the cubic term was assumed small, for reasons we will come to below). We can express these bounds in terms of a physical scales (inverse lengths) $\Lambda_{3}$ and $\Lambda_{4}$ that sets the size of the cubic and quartic powers of the Riemann operator in the action. These are related to our Wilson coefficients as
@@
$\Lambda^{-1}_{4}\lesssim\Lambda^{-1}_{4,0}\simeq 150\,\text{km}$ (4)
(i.e., we are using the notation $\Lambda_{4,0}$ as short-handed notation for the inverse scale of 150 km). This scale naturally emerges since black holes of mass $\sim 30\,M_{\text{sun}}$ have a Schwarzschild radius of $\approx 90$ km, which is of this order. Although the cubic terms were not directly included in the analysis, a similar bound is expected for them also, i.e., $\Lambda_{3}^{-1}\lesssim\Lambda^{-1}_{3,0}\sim 150\,\text{km}$.
It is important to note that such a bound is stronger than current bounds from the solar system. The reason for this is as follows: even though solar system bounds have more precision, they involve much weaker curvatures. Furthermore, if we go to scales below this; for example down to several micro-meters, the Cavendish type tests of gravity are still less stringent, as again the curvatures produced by laboratory masses are extremely tiny. So at the level of classical field theory, these bounds from GWs of merging black holes provide the tightest known direct constraints on the coefficients of these higher dimension operators.
@@ 2.3 Constraints from Causality
As has been known for the past few years Camanho:2014apa , if one only includes $(\text{Riem})^{3}$ in the effective action for gravitation, and no other higher dimension operators, then there is superluminality. However, when the quartic operators $(\text{Riem})^{4}$ are included, then non-zero cubic terms are allowed, so long as an inequality between the cubic and quartic coefficients (which will be explained more in-depth in section 2.4), is satisfied (see Caron-Huot:2022ugt and Bern:2021ppb ). The inequality is
@@ 2.4 Higher Derivative Operators from Integrating Out Massive Particles
The above higher derivative operators ultimately arise from whatever is the underlying UV completion of gravity. However, since these Wilson coefficients flow with scale, we do not need to know the full UV completion details, such as string theory. Instead, we can imagine that there are many massive, but somewhat light, particles that we integrate out, and we generate these operators in the infrared. As a concrete example we shall focus on integrating out $N$-minimally coupled massive scalars $\phi_{i}$. The starting action is
$S=\uprightintop\displaylimits d^{4}x\sqrt{-g}\biggm($ $-\Lambda-\frac{2R}{\kappa^{2}}+\sum_{i}\frac{1}{2}(\partial\phi_{i})^{2}-\sum_{i}\frac{1}{2}m_{i}^{2}\phi_{i}^{2}+\mathcal{L}_{\text{matter}}+\cdots\biggm)$ (9)
@@
These coefficients have a summation over the index $i=1,\ldots,N$ for each of the particle species that can be in the loop. The coefficients change depending on the spin of the particle in the loop. For our case, we only quote the results for s=0s=0.222For the case of non-zero spin, one finds that $\alpha_{3}=0$ to this order. In this case, one can still proceed, but one should use Eq. (7) rather then Eq. (5), leading to similar results to those found in this section, up to logarithmic correction factors. The coefficients $\tilde{\alpha}_{4}$ and $\tilde{\alpha}_{3}$ will not be essential for us here. From (5), we can find a relation between $m_{i}$ and $N$ given the coefficients in (10) given as
@@
The numerical prefactors $(75/11)^{1/6}$ and $(15/4)^{1/4}$ are both approximately 1.4.
2.4.1 All Masses Equal
If there is only one type of massive particle being integrated out ($m=m_{\text{min}}=m_{1}=\ldots m_{N}$) this becomes
$\Lambda_{3}\geq 1.4\,m_{\text{min}},\quad\quad\Lambda_{4}\geq 1.4\,m_{\text{min}}.$ (14)
As we discuss in the next subsection, the prefactor in the above expression for $\Lambda_{4}$ and $\Lambda_{3}$ is larger if there are different masses. So the best case scenario for phenomenological purposes then is that all the masses are equal $m=m_{\text{min}}=m_{1}=\ldots=m_{N}$. The more general cases of a distribution of masses will be analyzed in the next subsection.
We can readily invert the above logic to bound $m_{\text{min}}$. If we use the observational bound on the EFT of Ref. Sennett:2019bpc of $\Lambda_{4,0}\simeq(150\text{km})^{-1}$ as a useful reference, then we obtain a bound on the mass of the lightest particle of
$m\leq m_{\text{min}}=9.5\times 10^{-13}\,\text{eV}(\Lambda_{4}/\Lambda_{4,0})$ (15)
@@ 2.4.1 All Masses Equal
with equality holding when we saturate the theoretical bound $m=m_{c}$. Although this is an extremely large $N$, it is important to note that such species are assumed to be in a hidden sector and not directly coupled to the Standard Model at tree level. Their only direct couplings to the Standard Model occur via quantum loops as we now explore and learn their possible observational consequences. One might expect that such a huge number of particles would already have been easily visible at colliders, for example from $e^{+}e^{-}$ annihilation into these scalars via the graviton in the s-channel; but it is important to recall that the proposal that is presented in Ref. Sennett:2019bpc is that of the “soft UV completion", meaning that one should only use the theory at very low energies.
2.4.2 Distribution of Masses
If there are species of particles with different masses, then it is simple to see that the factor involving sums of powers of masses obeys
$P\equiv\frac{\left(\sum_{i}(m_{\text{min}}/m_{i})^{2}\right)}{\left(\sum_{i}(m_{\text{min}}/m_{i})^{4}\right)}\geq 1$ (19)
since there is a fourth power in the denominator and only a second power in the numerator of terms that are each $m_{\text{min}}/m_{i}\leq 1$ from the definition of $m_{\text{min}}$. Here equality only occurs when all the masses are equal, but otherwise this factor $P$ is larger than unity. In the latter case, the above inequalities (13) are even sharper, i.e., the energy scales controlling the higher order operators in the EFT must be even larger than $1.4\,m_{\text{min}}$ the lightest mass being integrated out. Hence if the masses being integrated out are somewhat heavy, then these scales are correspondingly microscopic and one cannot use a large $N$ to alter this conclusion.
Let us now begin an exploration into this more general case of a distribution of masses, rather than the above special case of $m=m_{\text{min}}=m_{1}=\dots=m_{N}$.
Let us define a function $\rho(m)$, the density of particle masses per unit mass. The previous case in which all particles have the same mass corresponds to a delta-function
$\rho_{d}(m)=N\,\delta(m-m_{\text{min}}).\,\,\,\,\,\mbox{(Delta)}$ (20)
However, we now consider a spread out distribution. For this purpose let us consider the following pair of representative examples:
$\rho_{p}(m)={N\,m_{\text{min}}\over m^{2}},\,\,\,m\geq m_{\text{min}},\,\,\,\,\,\mbox{(Power Law)}$ (21)
$\rho_{e}(m)={N\over f}\exp\left(-(m-m_{\text{min}})/f\right),\,\,\,m\geq m_{\text{min}},\,\,\,\,\,\mbox{(Exponential)}$ (22)
with $\rho=0$ for $m<m_{\text{min}}$ in both cases. Note that we have normalized the distribution such that $\uprightintop\displaylimits dm\,\rho(m)=N$ in all cases. In the exponential distribution, the parameter $f$ controls how rapidly the exponental falls off at high masses. For small $f\ll m_{\text{min}}$, this approaches a delta-function. While for large $f\gg m_{\text{min}}$, the distribution is very broad.
Using the above distributions, the above prefactor $P$ can be evaluated. For the fixed mass case (delta function distribution), one has $P_{d}=1$, as remarked above. However, for these more general distributions, we obtain $P>1$. For smooth distributions, we define $P$ as
$P={\uprightintop\displaylimits dm\>\rho(m)(m_{\text{min}}/m)^{2}\over\uprightintop\displaylimits dm\>\rho(m)(m_{\text{min}}/m)^{4}}$ (23)
The values are found to be
$P_{p}={5\over 3},\,\,\,\,\,\mbox{(Power Law)}$ (24)
$P_{e}=P_{e}(m_{\text{min}}/f),\,\,\,\,\,\mbox{(Exponential)}$ (25)
where $P_{e}$ is a dimensionless function of the ratio $m_{\text{min}}/f$; its explicit form is given in Appendix D. The function $P_{e}$ is a monotonic function of its argument, with limiting values:
$P_{e}\to 1,\,\,\,\,f\ll m_{\text{min}}$ (26)
$P_{e}\to 3,\,\,\,\,f\gg m_{\text{min}}$ (27)
This factor can then be inserted into Eq. (13). It is then raised to the power of $1/3$ for $\Lambda_{4}$ or the power of $1/4$ for $\Lambda_{3}$. This implies the corrections are only moderate. In particular we now have the bounds
$\Lambda_{4}\geq 1.7\,m_{\text{min}},\,\,\,\,\,\,\Lambda_{3}\geq 1.6\,m_{\text{min}}\,\,\,\,\,\,(\mbox{Power Law})$ (28)
$\Lambda_{4}\geq 2.0\,m_{\text{min}},\,\,\,\,\,\,\Lambda_{3}\geq 1.8\,m_{\text{min}}\,\,\,\,\,\,(\mbox{Exponential with}\,\,f\gg m_{\text{min}})$ (29)
In the exponential case we have reported on the $f\gg m_{\text{min}}$ limit. In the opposite limit, $f\ll m_{\text{min}}$ the distribution is that of a delta-function, so it just recovers the 1.4 prefactors of Eq. (14).
@@ 3.1 Massive Scalar Amplitude
$\Pi_{\lambda\xi\mu\nu,1}(q)$ $=-\frac{1}{2}\mu^{4-d}\uprightintop\displaylimits\frac{d^{d}l}{(2\pi)^{d}}\frac{\tau_{\lambda\xi}^{(1)}(l+q,l;m)\tau_{\mu\nu}^{(1)}(l,l+q;m)}{(l^{2}-m^{2})((l+q)^{2}-m^{2})}$
$=-\frac{1}{2}\mu^{4-d}\uprightintop\displaylimits\frac{d^{d}p}{(2\pi)^{d}}\frac{\tau_{\lambda\xi}^{(1)}(p+q,p;m)\tau_{\mu\nu}^{(1)}(p,p+q;m)}{(p^{2}-\Delta)^{2}}$ (33)
where $\Delta=m^{2}-x(1-x)q^{2}$ and $\mu$ is the scale at which we define the couplings, and we have suppressed the $i0^{+}$ in the Feynman propagator for the sake of brevity. The 1PI changes for equation (3.1) as
$\Pi_{\lambda\xi\mu\nu,2}(q)=\mu^{4-d}i\uprightintop\displaylimits\frac{d^{d}l}{(2\pi)^{d}}\frac{\tau^{(2)}_{\lambda\xi\mu\nu}{(l;m)}}{l^{2}-m^{2}}$ (34)
@@ 3.1.1 One-Loop Diagrams
$A_{0}(m^{2})=\frac{\mu^{4-d}}{(2\pi)^{d}}\uprightintop\displaylimits\frac{d^{d}p}{i\pi}\frac{1}{p^{2}-m^{2}}$ (37)
@@
$B_{0}(p^{2},m^{2},m^{2})=$ $\frac{\mu^{4-d}}{(2\pi)^{d}}\uprightintop\displaylimits\frac{d^{d}q}{i\pi^{2}}\frac{1}{(q^{2}-m^{2})((q+p)^{2}-m^{2})}$ (39)